### v2

# Suri: Multi-constraint Instruction Following for Long-form Text Generation

Abstract
Existing research on instruction following largely focuses on tasks with simple instructions and short responses. In this work, we explore multi-constraint instruction following for generating long-form text. We create Suri, a dataset with 20K human-written long-form texts paired with LLM-generated backtranslated instructions that contain multiple complex constraints. Because of prohibitive challenges associated with collecting human preference judgments on long-form texts, preference-tuning algorithms such as DPO are infeasible in our setting; thus, we propose Instructional ORPO (I-ORPO), an alignment method based on the ORPO algorithm. Instead of receiving negative feedback from dispreferred responses, I-ORPO obtains negative feedback from synthetically corrupted instructions generated by an LLM. Using Suri, we perform supervised and I-ORPO fine-tuning on Mistral-7b-Instruct-v0.2. The resulting models, Suri-SFT and Suri-I-ORPO, generate significantly longer texts ($\sim$5K tokens) than base models without significant quality deterioration. Our human evaluation shows that while both SFT and I-ORPO models satisfy most constraints, Suri-I-ORPO generations are generally preferred for their coherent and informative incorporation of the constraints. [^1] Code & Data are available at https://github.com/chtmp223/suri
Suri: Multi-constraint Instruction Following for
Long-form Text Generation
Chau Minh Pham  Simeng Sunthanks: Now at NVIDIA  Mohit Iyyer University of Massachusetts Amherst {ctpham,simengsun,miyyer}@cs.umass.edu
1 Introduction
Figure 1: Our work consists of two stages. First, we construct the Suri dataset using gold responses sampled from three existing datasets that include creative writing and open web text, along with backtranslated instruction $x_{w}$ and corrupted instruction $x_{l}$. Second, we fine-tune Mistral-7B-Instruct-v0.2 on Suri, resulting in two variations: Suri-I-ORPO (via I-ORPO) and Suri-SFT (via supervised fine-tuning).
Improving the instruction-following abilities of modern large language models (LLMs) is critical to increasing their effectiveness and generalizability for many practical applications. However, most existing instruction-following datasets (e.g., Alpaca) contain only simple instructions that can be solved by short model generations (Taori et al., 2023; Conover et al., 2023; Köpf et al., 2023). What about tasks with complex, multi-constraint instructions that can only be satisfied with long-form outputs (i.e., thousands of tokens), such as creating detailed technical reports or writing engaging fictional narratives?
We explore this question by conducting the first in-depth study of long-form instruction following with multi-constraint instructions. To facilitate our experiments, we create a new dataset, Suri, [^2] Suri is an alpaca breed known for its long, lustrous hair. using instruction backtranslation (Li et al., 2023; Köksal et al., 2023). This process involves feeding a human-written long-form text (e.g., chapters from a novel) into an LLM to generate instructions that could have been followed to create the text. The resulting dataset, Suri, consists of 20K texts paired with LLM-generated instructions, each containing $\approx$10 semantic and stylistic constraints (Figure 1).
How can we use Suri to improve an LLM’s long-form instruction following abilities? While supervised fine-tuning (SFT) has been quite effective for short-form datasets (Mishra et al., 2022; Wang et al., 2022; Sanh et al., 2022; Wei et al., 2022; Chung et al., 2022), we observe that fine-tuned Suri models often generate texts that are incoherent and fail to satisfy constraints towards the end in the instructions. Preference tuning methods such as DPO (Rafailov et al., 2023) and RLHF (Ouyang et al., 2022) are challenging to use in this setting due to difficulties and cost in obtaining preference judgments on long-form texts (Touvron et al., 2023; Xu et al., 2023c). Specifically, when annotating preferences for long texts, human annotators may struggle to determine if different sections of the text are faithful to the instructions while simultaneously considering multiple aspects of the text, such as coherence and informativeness.
Motivated by this, we devise an alignment method that relies on synthetically corrupted instructions. Specifically, we take the backtranslated instruction $x_{w}$ and corrupt its constraints using an LLM such that the gold response does not satisfy the corrupted constraints (for example, see $x_{l}$ in Figure 1). We then develop a variant of the Odds Ratio Preference Optimization objective (Hong et al., 2024, ORPO) to use these corrupted instructions as negative feedback. We refer to this alignment method as Instructional ORPO, or I-ORPO for short.
We conduct a series of automatic and human evaluations on generations from SFT and I-ORPO-tuned models to validate our method. Compared to the base model, Mistral-7b-Instruct-v0.2 (Jiang et al., 2023), both SFT and I-ORPO significantly increase the generation length from 1K to 5K tokens. Our fine-tuned models also improve the ability to differentiate between correct and corrupted instructions by at least 10% while maintaining low levels of $n$-gram repetitions in the text. We find that LLM judges, such as GPT-4o (OpenAI, 2024), Gemini-1.5-Pro (Team et al., 2024), and Claude-3.5-Sonnet (Anthropic, 2024), cannot reliably evaluate long-form responses, which makes human evaluation crucial for assessing the constraint-following capabilities of our generations. Annotators note that our fine-tuned models effectively follow given constraints, with I-ORPO being preferred to our SFT model for its ability to incorporate constraints coherently, informatively, and enjoyably.
2 The Suri Dataset
Category Dataset Size Domain Prompt Length Response Length
Writing Instructions ROCStory (Mostafazadeh et al., 2016) 50K Creative Writing 36 8
WritingPrompt (Fan et al., 2018) 300K Creative Writing 28 735
Instruction-following Alpaca (Taori et al., 2023) 52K General Q&As 13 44
Long-form Instruction-following LongForm-C (Köksal et al., 2023) 28K CommonCrawl, Wikipedia, StackExchange, Wikihow 149 298
LongAlpaca-16K (Chen et al., 2024b) 12K Science, Creative Writing, Wiki, General Q&As 5945 218
Scrolls (Shaham et al., 2022) 120K Legal, Science, Entertainment 33506 97
Multi-constraint Instructions Dolomites (Malaviya et al., 2024) 2K 25 Academic Fields 235 343
Multi-constraint, Long-form Instruction-following Suri (this work) 20K CommonCrawl, Creative Writing 347 4371
Table 1: Comparison of Suri with other single-turn datasets in terms of relevant data features, including writing instructions, instruction-following datasets, and constrained instructions. The data size as well as the average length of prompts and responses is either quoted from the original paper or calculated from publicly available subsets. Suri is the only dataset featuring both constrained instructions and long responses (> 4K tokens) specifically designed for text generation.
We focus on the task of long-form writing, both fictional and non-fictional, under multiple constraints. When using an LLM for a complex writing task, users might have many constraints in mind and expect lengthy, detailed responses in the form of books, blog posts, etc. This task is particularly challenging for current LLMs, which struggle with generating coherent long-form outputs Guan et al. (2021); Wang et al. (2023a), and this difficulty can be amplified when multiple constraints are involved. Recent instruction-following datasets have featured multi-constraint instructions (Xu et al., 2024; Malaviya et al., 2024) and long-form responses (Köksal et al., 2023; Chen et al., 2024b), but none has integrated these two elements (Table 1). We bridge this gap by creating Suri, which features complex instructions with multiple constraints and lengthy gold responses (2-5K words, about 3-6K tokens).
We collect human-written English text samples, such as books, religious texts, and blog posts, to serve as gold responses ($y$). Since gathering human-written instructions for such lengthy responses is difficult and expensive, we turn to instruction backtranslation (Li et al., 2023; Köksal et al., 2023), in which an LLM is provided with a human-written text (e.g., a short story) and prompted to generate instructions ($x_{w}$) that could have been followed to create that text. We further corrupt the constraints in $x_{w}$ to obtain synthetically corrupted instructions ($x_{l}$) for our I-ORPO alignment method. In total, Suri contains 20K single-turn examples, each consisting of a backtranslated instruction $x_{w}$, corrupted instruction $x_{l}$, and a human-written response $y$. In this section, we detail our approach to selecting high-quality text samples (§$\S$2.1) and creating backtranslated instructions (§$\S$2.2). We also validate our generated instructions (§$\S$2.3) and analyze the resulting dataset (§$\S$2.4).
2.1 Collecting Responses
Obtaining long-form gold responses $y$ through crowdsourcing or hiring experts requires significant cost and effort. As an alternative, we sample human-written texts in equal proportions from three existing datasets: ChapterBreak (Sun et al., 2022), Books3 (Presser, 2020; Gao et al., 2020), and RedPajama-Data-v2 (Computer, 2023). We truncate the sampled texts to between 2,048 and 5,024 words, making them significantly longer than those in existing instruction-following datasets (Table 1). The final Suri dataset is divided into training, validation, and test sets in a 10K/5K/5K split.
ChapterBreak
ChapterBreak (AO3 split) contains 7,355 fanfiction stories on Archive of Our Own (AO3), of which 6,656 texts are sampled for Suri. We merge the individual chapters from the cleaned text into a single document.
Books3
Books3 contains 197K books on Bibliotik, [^3] Due to copyright concerns, we only release the titles and IDs of the sampled data from this dataset. We provide a Python script to extract and clean the text so that users with access to Books3 can recreate the samples included in Suri. of which 6,698 texts are sampled. We use regular expressions to filter out irrelevant metadata, such as tables of contents and acknowledgments.
RedPajama-Data-v2
RedPajama contains over 100 billion documents from 84 CommonCrawl dumps. Unlike ChapterBreak and Books3, which consist primarily of books and literary narratives, RedPajama captures the style of everyday writing with informal textual content such as blog posts, obituaries, and more. We apply a set of quality filters (see Appendix A) on the 2023-06 and 2023-14 snapshots to obtain a subset of $\approx 300$K high-quality, non-duplicated documents written in English, from which 6,646 texts are sampled.
2.2 Creating Instructions via Backtranslation
Suri includes backtranslated instructions ($x_{w}$) and corrupted instructions ($x_{l}$). To create $x_{l}$, constraints from $x_{w}$ are minimally edited to be partially violated while still faithful to the overall main goal of the instruction. These corrupted instructions $x_{l}$, along with $x_{w}$ and $y$, serve as inputs for our I-ORPO preference tuning experiments.
Backtranslating Instructions
Our extracted gold responses do not come with accompanying instructions. Gathering these instructions can be costly and time-consuming, as annotators have to synthesize the instructions from long texts. Therefore, we use instruction backtranslation (Li et al., 2023; Köksal et al., 2023) to generate the missing instructions. Specifically, we prompt GPT-4-turbo [^4] GPT-4-turbo refers to gpt-4-0215-preview. Experiment done using temperature=0.6 and top_p=0.9. with a gold response $y$ to generate a corresponding instruction $x_{w}$ that contains a main goal, which summarizes the content of the text, and a list of $\approx$10 constraints (Table 9). These constraints can focus on stylistic elements (how something is communicated through tone, language, sentence structure), semantic elements (what topics, meanings, and concepts are included), or a combination of both. Constraints can also be broad, applying to large portions of the text, or specific, addressing elements that occur only once. The result is a set of highly detailed, multi-constraint instructions that cover different parts of the text ($x_{w}$ in Figure 1).
Corrupting Instructions
We want to use Suri in our alignment experiments, which traditionally rely on preference judgments (e.g., labeled $y_{w}$ and $y_{l}$ pairs). However, obtaining these judgments for long-form outputs is challenging due to the many competing aspects to consider (e.g., faithfulness to instructions, overall coherence, etc.). Therefore, we focus on learning from a corrupted instruction $x_{l}$ instead of from $y_{l}$. To create $x_{l}$, we prompt GPT-4-turbo [^5] Experiment done using model=gpt-4-0125-preview, temperature=0.0, top_p=0.0 to ensure deterministic results. to minimally edit each constraint in $x_{w}$ while preserving the original main goal (Table 10). The resulting instructions average 386 tokens, closely matching the average length of gold instructions at approximately 411 tokens. [^6] The first author manually reviews a random subset of 50 corrupted claims by comparing them to their corresponding original versions, and confirms that the corruptions are minimal and effective in invalidating the original constraints.
2.3 Validating Instructions
To validate whether the backtranslated instructions faithfully represent the original text, we conduct a human evaluation on a sample of 30 $(x_{w},y)$ pairs. Three Upwork [^7] See Appendix F for recruitment and compensation details. annotators are asked to read through the $(x_{w},y)$ pairs, highlight all text spans in the response that support the given constraints, and determine if the response supports the instruction (Figure 6).
Our findings indicate that, on average, about 87% of the constraints are fully satisfied, with the remaining constraints being partially satisfied (see Appendix F for agreement statistics). We conclude that the backtranslated instructions are generally faithful to the original text.
2.4 Instruction Diversity
Instructions in Suri focus primarily on long-form text generation, particularly crafting narratives or articles. Therefore, the key element that introduces diversity across these instructions is the accompanying list of constraints. Here, we measure the proportion of constraints being broad/specific or focusing on semantic/stylistic elements. We prompt Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), [^8] Experiment done using greedy decoding. The first author manually verifies an output subset. to assign each constraint to the applicable category. We find that semantic constraints account for more than half of each instruction, followed by mixed constraints (Figure 2). Broad constraints, on the other hand, make up 56% of the total constraints. Overall, the distribution of constraint types is relatively balanced, with a stronger emphasis on broad and semantic constraints.
Figure 2: Average percentage of different constraint types within each instruction. The left figure categorizes the constraints based on their content, and the right figure refers to constraint scopes.
3 Aligning language models with Suri
Our goal is to assess whether Suri helps improve the instruction-following capabilities of Mistral-7B-Instruct-v0.2 for long-form text generation.
We experiment with two methods of fine-tuning Mistral-7B-Instruct-v0.2 on Suri: supervised fine-tuning (SFT) using $(x_{w},y)$ pairs and a modified ORPO alignment (Hong et al., 2024) using $(x_{w},x_{l},y)$ triplets. We emphasize that preference judgments are difficult to obtain for long-form responses due to numerous aspects of the text that must be considered with respect to the instructions. Therefore, we perform model alignment with correct instruction $x_{w}$ and corrupted instruction $x_{l}$ instead. Full details on fine-tuning libraries, hardware configurations, and hyperparameters can be found in Appendix C.
Suri-I-ORPO
Odds Ratio Preference Optimization (ORPO) Hong et al. (2024) combines SFT and preference alignment by incorporating a log odds ratio term into the negative log-likelihood loss. We choose ORPO due to its simplicity, competitive performance with other preference tuning algorithms and the ease with which we can modify for our setting.
The original algorithm learns from preference judgments, requiring access to chosen and rejected responses in the $(x,y_{w},y_{l})$ format. Since our dataset contains gold and corrupted instructions instead, we modify ORPO so that the algorithm accepts $(x_{w},x_{l},y)$. We refer to this modified method as Instructional Odds Ratio Preference Optimization (I-ORPO), where the modified loss is calculated as:
$\mathcal{L}_{\text{I-ORPO}}=\mathbb{E}_{(x_{w},x_{l},y)}\left[\mathcal{L}_{\text{SFT}}+\lambda\cdot\mathcal{L}_{\text{I-OR}}\right]$ (1)
where
$\mathcal{L}_{\text{I-OR}}=-\log\sigma\left(\log\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}\right)$ (2)
$\textbf{odds}_{\theta}(y|x)=\frac{P_{\theta}(y|x)}{1-P_{\theta}(y|x)}$ (3)
In the original ORPO formulation, the model is learning if the log probability of $P_{\theta}(y_{w}|x)$, denoted $\text{logps}(y_{w}|x)$, increases and log probability of $P_{\theta}(y_{l}|x)$, denoted $\text{logps}(y_{l}|x_{w})$, decreases after a number of training steps, resulting in the log odds ratio increasing. In I-ORPO, the same $y$ is used for both instruction types. Therefore, the model is learning if the log probabilities $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$ diverge while $\text{logps}(x_{w})$ and $\text{logps}(x_{l})$
remain stable. We observe this trend in Figure 3. Loss derivation and analysis are in Appendix D.
Figure 3: ORPO training curve. Left figure documents the log probability of the chosen and rejected prompts over 3 epochs. Right figure shows the log probability of the response given the chosen and rejected prompts over 3 epochs. A divergence between $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$ is observed after 0.5 training epoch.
We perform I-ORPO fine-tuning with LoRA on Mistral-7B-Instruct-v0.2 for two epochs, using a learning rate of 5e-5, $\lambda$ of 0.4, and a LoRA rank and alpha of 16. We do not observe signs of the model learning with full-model tuning, so we choose to use LoRA fine-tuning instead. To minimize noise and improve the model’s ability to distinguish between gold and corrupted instructions, we include a single constraint in each instruction, $x_{w}$ and $x_{l}$.
Suri-SFT
We perform LoRA supervised fine-tuning Hu et al. (2021) on Mistral-7B-Instruct-v0.2 for two epochs using a learning rate of 5e-5, with a LoRA rank and alpha of 16. For each instruction $x_{w}$, we include a varying number of constraints to expose the model to different instruction formats. We do not use full-model tuning to match the I-ORPO training setting.
For both Suri-SFT and Suri-I-ORPO, we set the number of epochs to 2, which is determined by a manual inspection of model generations at each epoch checkpoint. We observe that as the number of epochs increases from 1 to 5, both the response length and the number of 5-gram repetitions increase, indicating a trade-off between response length and repetition. After reviewing 30 generations at each epoch, we select the configuration that produces the most coherent responses and minimal repetition. Based on these criteria, the optimal number of epochs is 2, which balances the trade-off between response length and response quality.
4 Automatic Evaluation
Our automatic assessment demonstrates that both Suri-I-ORPO and Suri-SFT increase the length of the generated texts while maintaining a reasonable level of repetition. Compared to baseline models, Suri-I-ORPO is more likely to assign higher log probabilities to tokens in the response given the correct instruction than the corrupted instruction.
4.1 Suri-I-ORPO and Suri-SFT generate substantially longer text.
We measure the average number of tokens [^9] Measured using tiktoken package (https://github.com/openai/tiktoken) with “o200k_base” encoding. in generations from our fine-tuned models (Suri-I-ORPO and Suri-SFT) and compare them to baseline models, including Mistral-7B-Instruct-v0.2, Llama-3-8B-Instruct (AI@Meta, 2024), and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). For faster inference, we use vLLM (Kwon et al., 2023) to generate outputs from the backtranslated instruction $x_{w}$. [^10] Experiment done using greedy decoding, max_token=10K. Inference prompts specify that 5K tokens should be generated. Proprietary models like GPT-4 and Claude are excluded due to their maximum generation output limit of 4,096 tokens, [^11] Claude documentation; OpenAI documentation whereas open-weight models allow for outputs of arbitrary maximum length.
Figure 4: Average number of tokens in generations from baseline open-source models (Llama-3-8B-Instruct, Mixtral-8x7B-Instruct-v0.1, Mistral-7B-Instruct-v0.2) and our fine-tuned models (Suri-I-ORPO, Suri-SFT).
Our fine-tuned models, Suri-SFT and Suri-I-ORPO, generate significantly longer outputs compared to the open-weight baselines, with an average of approximately 4,800 and 5,100 tokens per generation, respectively (Figure 4). These lengths exceed the maximum generation capacity of proprietary models, which is limited to around 4,096 tokens. Among the baselines, Mixtral produces the longest generations, averaging over 1,500 tokens, while Mistral-Instruct generates the shortest outputs, around 1,100 tokens per generation.
4.2 Suri-I-ORPO and Suri-SFT do not degenerate into repetitions at longer sequences.
We analyze the presence of repetitions in model generations. Since LLMs often degrade into repetitions over longer sequences, this measurement helps us identify when and how the model starts producing repetitive content. Previous work (Li et al., 2016; See et al., 2019) measures unigram, bigram, and trigram repetitions. However, we are interested in sentence-level repetitions, such as when the same phrase is repeated in a dialogue at the start of each sentence. Therefore, we measure 5- and 10-gram repetitions to capture these higher-level patterns. We count a repetition when a specific $n$-gram appears at least three times in the text.
I-ORPO SFT Mistral-Instruct Llama-Instruct Mixtral-Instruct
5-gram 24% 29% 12% 26% 31%
10-gram 3% 3% 1% 2% 5%
Table 2: Percentage of generations containing $n$-gram repetitions out of 5K generations from the test set (rounded to the nearest whole number).
Figure 5: Average percentage of 5-gram repetitions before and after 2,048 tokens in each generation from I-ORPO and SFT models.
Despite having the longest generations, Suri-I-ORPO and Suri-SFT maintain a low percentage of generations with $n$-gram repetitions (Table 2). Among the baseline models, Mistral-Instruct has the lowest percentage of generations with repetition, possibly because its generations are also the shortest. Surprisingly, Llama-Instruct and Mixtral-Instruct, with their short generations, possess a greater proportion of generations with $n$-gram repetitions compared to our fine-tuned models.
We further examine the percentage of 5-gram repetitions, normalized by the length of each text, generated by our fine-tuned models. As shown in Figure 5, the percentage of 5-gram repetitions does not increase after 2,048 tokens, indicating that our fine-tuned models do not exhibit degradation in longer sequences.
4.3 I-ORPO improves ranking accuracy
To understand the capabilities of models to differentiate between correct and corrupted instructions, we evaluate ranking accuracy (See et al., 2019; Chen et al., 2024a). This involves measuring the percentage of cases in which the model assigns a higher probability to the gold response under the correct instruction than under the corrupted version. We calculate the sum of token log probabilities in the response given the previous tokens, denoted by $\text{logps}(y|x)$, and determine accuracy based on the proportion of times when $\text{logps}(y|x_{w})>\text{logps}(y|x_{l})$. A higher accuracy indicates that the model is more sensitive to the instructions and can determine which instruction is the correct instruction for the given response.
We use Hugging Face’s Transformers Wolf et al. (2020) to access the probability distribution over vocabulary and measure the impact of instruction specificity on ranking accuracy across five different settings, which are defined by the number of all constraints included (M constraints in total) and the number of those included constraints that are corrupted: (M,M), (M,M/2), (M,1), (M/2,M/2), (1,1). For example, in the (M, M/2) setting, both instructions include all constraints, but only half of the constraints are violated.
Instruction Specificity I-ORPO SFT Mistral-Instruct Llama-Instruct Mixtral-Instruct
(M,M) 100.0 99.8 90.6 65.7 66.5
(M,M/2) 100.0 99.2 92.1 57.5 60.4
(M,1) 98.3 91.0 90.4 47.7 55.2
(M/2,M/2) 99.9 97.8 79.7 60.0 57.4
(1,1) 98.4 81.2 62.5 50.9 48.5
Table 3: Ranking accuracy on the Suri test set across five levels of instruction specificity. Percentages are rounded to one decimal place.
Suri-I-ORPO shows at least a 10% improvement in ranking accuracy over the baseline Mistral-Instruct across all instruction specificity settings, with Suri-SFT following closely (Table 3). Mistral-Instruct remains a strong baseline, achieving the highest ranking accuracy among the three baseline models. In contrast, Llama-3-7b-Instruct and Mixtral-8x7b-Instruct perform the worst, trailing Suri-I-ORPO by up to 50%. We observe that settings with more constraints in the instruction, namely (M,M), (M,M/2), and (M,1), generally lead to better performance. This trend suggests that seeing more constraints helps the model better differentiate between correct and corrupted constraints.
4.4 LLM judges are unreliable for evaluating constraint satisfaction in long-form generation.
We experiment with using LLMs to evaluate whether texts generated by our fine-tuned models follow the given constraints. Specifically, we provide GPT-4o (OpenAI, 2024), Gemini-1.5-Pro (Team et al., 2024), and Claude-3.5-Sonnet (Anthropic, 2024) with a constraint and a generated text from our models and prompt it to determine whether the text fully satisfies, partially satisfies, or does not satisfy the constraint (Table 16). [^12] Experiment done using OpenAI API for GPT-4o and Vertex API for Claude-3.5-Sonnet and Gemini-1.5-Pro. Temperature is set to 0.0 and the maximum number of generated tokens is set to 4096 for all models. We then compare these results with judgments from three Upwork annotators on 30 texts generated by Suri-SFT on the test set (obtained using the same procedure as in Section 2.3). GPT-4o agrees with human annotators only 39% of the time, with a significant 16% disagreement between satisfaction and no satisfaction (Table 4). Claude-3.5-Sonnet and Gemini-1.5-Pro lagging significantly behind, with Claude agreeing with humans only 24% of the time, and Gemini 13% of the time. Notably, Gemini refuses to annotate 39% of the time, even when all safety filters have been disabled. We conclude that LLM judges do not align well with long-form human annotation, consistent with the findings of Xu et al. (2024) and Kim et al. (2024).
Category GPT-4o Claude-3.5-Sonnet Gemini-1.5-Pro
Agreement with human’s majority vote 39% 24% 13%
Partial satisfaction - No satisfaction 23% 6% 2%
Satisfaction - Partial satisfaction 22% 63% 1%
Satisfaction - No satisfaction 16% 7% 45%
Table 4: Types of agreement and disagreement between GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and human judges on 30 generations from Suri-SFT.
5 Human Evaluation
While our automatic assessments provide insights into the lexical information of the text, they do not capture its semantic content. Therefore, we conduct a human evaluation to determine if and how the constraints are satisfied by the outputs of Suri-SFT and Suri-I-ORPO. Human evaluation on 30 test set generations reveals that while both fine-tuned models satisfy constraints, Suri-I-ORPO is preferred by humans for its ability to seamlessly incorporate the constraints into the final outputs.
5.1 Suri-I-ORPO and Suri-SFT are effective at satisfying constraints.
Since GPT-4o judgments do not align with human annotations, we rely on human evaluation to determine how often Suri-I-ORPO and Suri-SFT follow the given constraints. This evaluation follows a similar setup as Section 2.3, where annotators assess whether each constraint is satisfied, partially satisfied, or not satisfied by the generations. Two Upwork annotators complete 30 tasks, each containing a generation with around ten constraints, totaling 321 constraints. The generations are lengthy, averaging 4,000 words, and complex, with constraints spread throughout the text. Annotators spend approximately 20-25 minutes on each annotation and are paid \$200 in total for the task.
On average, Suri-I-ORPO and Suri-SFT meet most of the included constraints, achieving satisfaction rates of 67-68% and partial satisfaction rates of 16-17% (Table 5). Both models have the same proportion of unsatisfied constraints, accounting for 16% of the total constraints. Annotators often note that narratives produced by Suri-SFT contain inconsistent plot events and sometimes leave the narrative incomplete, resulting in some final constraints not being met. We attribute this behavior to the fact that some of the gold responses are truncated to between 2,048 and 5,024 words, which might omit the end of the original narrative. On the other hand, Mistral-I-ORPO produces narratives with coherent endings but can sometimes be too verbose, making it difficult to determine whether some constraints are satisfied.
Suri-SFT Suri-I-ORPO
Satisfied 67% 68%
Partially Satisfied 17% 16%
Not Satisfied 16% 16%
Table 5: Average percentage of satisfied constraints in Suri-SFT and Suri-I-ORPO generations. Percentages are rounded to the nearest whole number.
5.2 Suri-I-ORPO are preferred over Suri-SFT for coherent and informative constraint satisfaction.
In this evaluation, we are interested in how our fine-tuned models satisfy constraints in Suri. We ask two annotators to compare text generations from Suri-SFT and Suri-I-ORPO with respect to a given constraint based on the following criteria:
  • Informativeness: Which generation provides more details about the constraint?
  • Coherence: Which generation effectively integrates the constraint with the rest of the text?
  • Readability/Enjoyability: Which text sample is easier to read overall?
The annotators also provide detailed justifications for their choices in each aspect of their judgments (see Figure 7).
Coherence Informativeness Enjoyability/Readability
72% 73% 67%
Table 6: Win rate of Suri-I-ORPO over Suri-SFT in terms of coherence, enjoyability, and informativeness.
Human annotators consistently prefer Suri-I-ORPO to Suri-SFT for about 60-70% of the time across all three categories: coherence, informativeness, and enjoyability. Annotators note that Suri-SFT often suffers from repetitive ideas, confusing plot points, and a lack of proper conclusions. In contrast, while Suri-I-ORPO texts occasionally exhibit inconsistencies, they generally read more naturally, include interesting details, and are devoid of the robotic structure or flowery language often found in other LLM generations.
6 Related Work
Instruction Following Datasets
Open-ended instruction tuning involves fine-tuning LLMs to follow user instructions and generate high-quality text (Wei et al., 2021; Askell et al., 2021; Ouyang et al., 2022; Liu et al., 2023; Rafailov et al., 2023). Single-turn instruction-following datasets can be constructed by manual annotation, where instruction-response pairs are curated by humans (Conover et al., 2023; Rajani et al., 2023; Zhou et al., 2024). Another approach is distillation from proprietary LLMs, which can be done via techniques like Self-instruct (Wang et al., 2023c) to augment responses for each instruction (Taori et al., 2023; Xu et al., 2023a, b), Instruction Backtranslation to generate instructions given gold responses (Köksal et al., 2023; Li et al., 2023), or leveraging metadata to generate both instructions and responses (Yin et al., 2023). While recent work has constructed instruction-following datasets with long-form responses (Xiong et al., 2023; Chen et al., 2024b; Bai et al., 2024) or multiple constraints (Xu et al., 2024; Zhou et al., 2023; Malaviya et al., 2024), no prior effort has explored combining these two elements in single-turn instructions (see Table 1). Suri is the first dataset to feature both complex instructions and long-form responses over 5k words.
Alignment
Aligning language models with instruction-following data is crucial for ensuring that they respond to user instructions in a helpful and harmless manner (Askell et al., 2021; Mishra et al., 2022; Sanh et al., 2022; Chung et al., 2022; Wang et al., 2023b). Popular preference tuning methods, such as RLHF, DPO, KTO, and ORPO Ouyang et al. (2022); Rafailov et al. (2023); Ethayarajh et al. (2024); Hong et al. (2024), achieve this by fine-tuning the models on human judgments of response quality (Kreutzer et al., 2018; Stiennon et al., 2022; Ziegler et al., 2020; Ramamurthy et al., 2023). However, collecting preferences for long-form responses is challenging due to the many competing aspects of the texts that need to be considered, such as instruction faithfulness and coherence (Xu et al., 2023c; Kim et al., 2024; Xu et al., 2024), which prompts us to experiment with preference tuning on correct and correct instructions.
7 Conclusion
In this work, we investigate the challenge of complex instruction following for generating long-form text. We introduce Suri, a dataset of long human-written responses accompanied by backtranslated and corrupted instructions. We demonstrate the effectiveness of Suri in improving the constraint-following capabilities of LLMs for long-form generation through supervised fine-tuning and I-ORPO. Human and automated evaluations show that our models generate high-quality, long-form responses while effectively satisfying constraints.
Limitations
Fine-tuning additional LLMs on Suri
While we demonstrate the effectiveness of Suri and I-ORPO on Mistral-7b-Instruct-v0.2, we have yet to experiment with fine-tuning other models on our dataset using I-ORPO.
Impact of surface features on I-ORPO
Even though I-ORPO works well on our dataset, we would like to explore how surface features, such as instruction length and the degree of information overlap between the instruction and response, affect its performance.
Impact of truncating gold responses
In our experiments, we truncate gold responses to lengths between 2,048 and 5,024 words to make fine-tuning more cost-effective and computationally efficient. However, our released code includes an option that allows users to recover the full response text, and thus bypass the truncated version if needed.
Ranking accuracy on out-of-domain datasets
We report the ranking accuracy on the Suritest set, where Suri-SFT and Suri-I-ORPO may have an advantage over the baseline models due to their fine-tuning on Suri.
Ethical Considerations
Our human evaluation receives approval from an institutional review board. All annotators (US-based, fluent in English) gave their informed consent and participated with an hourly compensation of \$16, which meets the minimum wage in our state. Scientific artifacts are implemented according to their intended usage.
Acknowledgements
We extend special gratitude to the Upwork annotators for their hard work and to the members of Unsloth, r/LocalLLaMA, and Together.ai community for helpful fine-tuning advice. We also thank Scott Niekum, Dzung Pham, and members of the UMass NLP lab for their insights on the project. This project was partially supported by awards IIS-2202506 and IIS-2312949 from the National Science Foundation (NSF) and an award from Open Philanthropy.
Appendix A Quality Filters for RedPajama-Data-v2
Upon initial examination, we observe a significant presence of news and religious text in the corpus. Therefore, in addition to the following quality filters, we also downsample news and religious articles by excluding any article containing a source domain on our blocklist (BernhardClemm, 2023) or more than 0.05% of words from a religious dictionary (The Association of Religion Data Archives, 2023) to ensure the diversity of the gold responses. Table 7 and 8 show the quality filters used in RedPajama-Data-v2.
Tags Values Descriptions Categories
ccnet_language_score $>0.65$ score of the language identification model CCNet
ccnet_perplexity $(35,350)$ perplexity of an LM trained on Wikipedia CCNet
rps_doc_books_importance $>0$ Given a bag of {1,2}-wordgram model trained on Books p, and a model trained on the source domain q, This is the logarithm of the ratio p(doc)/q(doc). ML Heuristics
rps_doc_curly_bracket 00 The ratio between the number of occurrences of ’{’ or ’}’ and the number of characters in the raw text. Some pages inadvertently contained code. Since the curly bracket, "{" appears in many programming languages (such as Javascript, widely used on the web) but not in natural text, we removed any pages that contained a curly bracket. Natural Language
rps_doc_frac_no_alph_words $0.3$ The fraction of words that contain no alphabetical character. Natural Language
rps_doc_lorem_ipsum 00 The ratio between the number of occurrences of ’lorem ipsum’ and the number of characters in the content after normalisation. Natural Language
rps_doc_unigram_entropy $\geq 3$ The entropy of the unigram distribution of the content. This measures the diversity of the content and is computed using sum(-x / total * log(x / total)) where the sum is taken over counts of unique words in the normalised content. Natural Language
Table 7: Quality Signals used to filter RedPajamas Dataset - Part 1
Tags Values Descriptions Categories
rps_doc_word_count $(2048,5024)$ The number of words in the content after normalisation. Natural Language
[…]
rps_doc_frac_chars_dupe_10grams 0.1 The fraction of characters in duplicate word ngrams. Repeti-tiveness
rps_doc_frac_chars_dupe_5grams $0.15$
rps_doc_frac_chars_dupe_6grams $0.14$
rps_doc_frac_chars_dupe_7grams $0.13$
rps_doc_frac_chars_dupe_8grams $0.12$
rps_doc_frac_chars_dupe_9grams $0.11$
rps_doc_frac_chars_top_2gram $0.2$
rps_doc_frac_chars_top_3gram $0.18$
rps_doc_frac_chars_top_4gram $0.16$
rps_doc_ldnoobw_words 00 The number of sequences of words that are contained in the List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words blocklist. The blocklist is obtained from https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words Sensitive / toxic content
rps_doc_ut1_blacklist 00 A categorical id corresponding to the list of categories of the domain of the document. Categories are obtained from the UT1 blacklist. The list is obtained from https://dsi.ut-capitole.fr/blacklists/: [’adults’, ’phishing’, ’dating’, ’gambling’, ’filehosting’, ’aggressif’, ’ddos’, ’mixed_adult’, ’chat’, ’arjel’] Sensitive / toxic content
Table 8: Quality Signals used to filter RedPajamas Dataset - Part 2
Appendix B Prompts
In this section, we show prompts to generate and analyze Suri in Table 9, 10, 11, 12. Table 16 shows the prompt used for our experiment with LLM judges.
Prompt: Instruction Backtranslation/Reverse-engineering
Assume the author of the provided text followed a detailed set of instructions to produce their work. Your task is to infer what those original instructions may have been by composing your own set of instructions that could recreate key aspects of the given text.
Your response must include: 1. An overarching instruction under the "Main Instruction" section that summarizes the goal of the instructions. 2. One bulleted list of specific constraints under the "Constraints" section that reflect the order of happenings/ideas in the original text. Constraints should focus on either stylistic elements (how something is communicated through tone, language, sentence structure), semantic elements (what topics, meanings, and concepts are included), or a combination of both. You should include specific elements from the text, but avoid using direct quotes. Aim for a fair balance of semantic, stylistic and mixed constraints. Examples of stylistic constraints are "incorporate humor when discussing serious topics" or "use short, choppy sentences for emphasis." Examples of semantic constraints are "describe a supportive mother and absent father" or "mention an impressionist painting with a leopard." Mixed constraints blend stylistic and semantic elements, like "discuss impressionist art with an enthusiastic tone." ### Document:
{text}
### Your response:
[…]
Prompt: Corrupt backtranslated instructions
You are given an instruction text that includes a main instruction and a list of constraints. Your task is to make minimal edits to violate each constraint. Your resulting constraints should be coherent with one another and also with the main instruction.
[Examples]
Main Instruction: Write a story on the life and death of Bob, who is a run-of-the-mill blue-collar worker in Texas, USA.
Constraints: Use a first-person perspective that centers on the protagonist’s perspective. $\rightarrow$ Use a third-person perspective that ensures a broad and neutral view of the narrative. Include cliffhangers at the end of each chapter to encourage readers to continue reading. $\rightarrow$ Do not include cliffhangers at the end of each chapter to encourage smooth readings.
[Provided Instruction]
{instructions}
When modifying the constraints, keep the following in mind: 1. Ensure that your resulting constraints are coherent with one another and also with the main instruction. However, the original and modified constraints should be mutually exclusive and difficult to achieve simultaneously. 2. Modify every constraint, but leave the main instruction unchanged. 3. Your response should contain the original main instruction, followed by each original constraint and your minimally modified version. Format each constraint as: Original constraint $\rightarrow$ Your modified constraint. [Your response]
Table 10: Prompt used to violate backtranslated instructions. The placeholder {instructions} are replaced with instructions that are produced with Prompt 9.
Prompt: Assign constraint type (semantic, stylistic, mixed) to each constraint
You are a helpful assistant. You are given a constraint that you need to determine if it is a stylistic, semantic, or mixed constraint. Stylistic constraint emphasizes stylistic elements (how something is communicated through tone, language, sentence structure). Stylistic constraints focus on semantic elements (what topics, meanings, and concepts are included). Mixed constraints include both stylistic and semantic elements.
### Examples:
Constraint: Incorporate humor when discussing the morbid, gut-wrenching scene of the protagonist’s death. Use short, choppy sentences to create a sense of urgency and panic.
Your response: Stylistic
Constraint: The story must end with the protagonist’s death in a car accident.
Your response: Semantic
Constraint: Using a first-person perspective, write a story on the life and death of Bob, a blue-collar worker in Texas, USA.
Your response: Mixed
Constraint: Include cliffhangers at the end of each chapter to encourage readers to continue reading.
Your response: Stylistic
### Constraints:
Constraint: {constraint}
### Your response:
Table 11: Prompt to assign constraint type (semantic, stylistic, mixed) to each constraint. The placeholder {constraint} will be replaced with a single constraint in each backtranslated instruction.
Prompt: Assign constraint scope (broad, specific) to each constraint
You are a helpful assistant. You are given a constraint that you need to determine if it is a specific or broad constraint. Specific constraints focus on an element that can be found in a specific part of the text. Broad constraints focus on an element that can be found throughout the text.
### Examples:
Constraint: Throughout the narrative, use a first-person perspective that centers on the protagonist’s perspective.
Your response: Broad
Constraint: Include cliffhangers at the end of the first chapter to encourage readers to continue reading.
Your response: Specific
Constraint: Introduce a new character in the middle of the story to add depth to the narrative.
Your response: Specific
Constraint: Include cliffhangers at the end of each chapter to encourage readers to continue reading.
Your response: Broad
### Constraints:
Constraint: { $x_{w}$ constraint}
### Your response:
Table 12: Prompt to assign constraint scope (broad/specific) to each constraint. The placeholder {$x_{w}$ constraint} is replaced with a single constraint from each backtranslated instruction.
Appendix C Modeling Experiment Details
All experiments are done using Flash-Attention 2 (Dao, 2024), DeepSpeed ZeRO 3 (Rasley et al., 2020), PEFT (Mangrulkar et al., 2022), TRL library (von Werra et al., 2020), and Alignment Handbook (Tunstall et al., 2023). Chat templates are as follows:
<|user|>
{Instruction}</s>
<|assistant|>
{Response}</s>
The training configurations (Table 13) are mostly similar for SFT and ORPO. We vary the learning rate (5e-4 to 5e-7), optimizer (8-bit vs. 32-bit), LoRA rank, and alpha (8 to 64), but none of these hyperparameters results in better generations.
Configurations Values
Hardware (Training and Inference) 4xA100s
Tracking wandb
lora_r 16
lora_alpha 16
lora_dropout 0.05
beta (for ORPO only) 0.4
gradient_accumulation_steps 1
gradient_checkpointing True
learning_rate 5.0e-5
lr_scheduler_type cosine
max_length 15024
max_completion_length 15000
max_prompt_length 5000
num_train_epochs 2
optim adamw_torch
per_device_train_batch_size 1
Table 13: Training details for SFT and ORPO
Appendix D I-ORPO Loss Derivation
The derivation of $\mathcal{L_{\text{I-OR}}}$ closely resembles that of the original ORPO loss, with $d=(x_{w},x_{l},y)\sim D$.
$\nabla_{\theta}\mathcal{L_{\text{I-OR}}}=\delta(d)\cdot h(d)$ (4)
$\delta(d)$ $=\left(1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}\right)^{-1}$ (5)
$h(d)$ $=\frac{\nabla_{\theta}\log P_{\theta}(y|x_{w})}{1-P_{\theta}(y|x_{w})}-\frac{\nabla_{\theta}\log P_{\theta}(y|x_{l})}{1-P_{\theta}(y|x_{l})}$ (6)
The gradient of $\mathcal{L_{\text{I-OR}}}$ is the product of two terms: $\delta(d)$, which regulates the strength of parameter updates, and $h(d)$, which widens the contrast between $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$. Specifically, as the odds ratio increases, $\delta(d)$ converges to 0. On the other hand, $h(d)$ has two gradients: $\nabla_{\theta}\log P_{\theta}(y|x_{w})$, which minimizes $\log P_{\theta}(y|x_{w})$, and $\nabla_{\theta}\log P_{\theta}(y|x_{l})$, which maximizes $\log P_{\theta}(y|x_{l})$. Additionally, $1-P_{\theta}(y|x_{w})$ accelerates the update in the direction that maximizes $P_{\theta}(y|x_{w})$. Following ORPO Hong et al. (2024), suppose that $g(x_{w},x_{l},y)=\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}$, we derive the loss as in 22.
$\nabla_{\theta}\mathcal{L}_{I-OR}$ $=\nabla_{\theta}\log\sigma\left(\log\left(\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}y|x_{l})}\right)\right)$ (7)
$=\frac{1}{\sigma(\log g(x_{w},x_{l},y))}\cdot\nabla_{\theta}\sigma(\log g(x_{w},x_{l},y))$ (8)
$=\frac{1}{\sigma(\log g(x_{w},x_{l},y))}\cdot\sigma(\log g(x_{w},x_{l},y))(1-\sigma(\log g(x_{w},x_{l},y)))\nabla_{\theta}\log g(x_{w},x_{l},y)$ (9)
$=(1-\sigma(\log g(x_{w},x_{l},y)))\cdot\nabla_{\theta}\log g(x_{w},x_{l},y)$ (10)
$=\sigma(-\log g(x_{w},x_{l},y))\cdot\nabla_{\theta}\log g(x_{w},x_{l},y)$ (11)
$=\left(1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}\right)^{-1}\cdot\nabla_{\theta}\log\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}$ (12)
$=\left(1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}\right)^{-1}\cdot\nabla_{\theta}\log\left(\frac{P(y|x_{w})}{1-P(y|x_{w})}\frac{1-P(y|x_{l})}{P(y|x_{l})}\right)$ (13)
$\nabla\log\left(\frac{P(y|x_{w})}{1-P(y|x_{w})}\frac{1-P(y|x_{l})}{P(y|x_{l})}\right)$ can be rewritten as:
$=\nabla_{\theta}\log\left(\frac{P(y|x_{w})}{P(y|x_{l})}\frac{1-P(y|x_{l})}{1-P(y|x_{w})}\right)$ (14)
$=\nabla_{\theta}\log\left(\frac{P(y|x_{w})}{P(y|x_{l})}\frac{1-P(y|x_{l})}{1-P(y|x_{w})}\right)$ (15)
$=\nabla_{\theta}\log\frac{P(y|x_{w})}{P(y|x_{l})}-(\nabla_{\theta}\log(1-P_{\theta}(y|x_{w}))-\nabla_{\theta}\log(1-P_{\theta}(y|x_{l})))$ (16)
$=\nabla_{\theta}\log\frac{P(y|x_{w})}{P(y|x_{l})}-\left(\frac{\nabla_{\theta}(1-P_{\theta}(y|x_{w}))}{1-P_{\theta}(y|x_{w})}-\frac{\nabla_{\theta}(1-P_{\theta}(y|x_{l}))}{1-P_{\theta}(y|x_{l})}\right)$ (17)
$=\nabla_{\theta}\log\frac{P(y|x_{w})}{P(y|x_{l})}-\left(\frac{-\nabla_{\theta}(P_{\theta}(y|x_{w}))}{1-P_{\theta}(y|x_{w})}-\frac{-\nabla_{\theta}(P_{\theta}(y|x_{l}))}{1-P_{\theta}(y|x_{l})}\right)$ (18)
$=\nabla_{\theta}\log\frac{P(y|x_{w})}{P(y|x_{l})}-\left(\frac{-P_{\theta}(y|x_{w})\nabla_{\theta}\log P_{\theta}(y|x_{w})}{1-P_{\theta}(y|x_{w})}-\frac{-P_{\theta}(y|x_{l})\nabla_{\theta}\log P_{\theta}(y|x_{l})}{1-P_{\theta}(y|x_{l})}\right)$ (19)
$=\nabla_{\theta}\log\frac{P(y|x_{w})}{P(y|x_{l})}-\left(-\textbf{odds}_{\theta}(y|x_{w})\cdot\nabla_{\theta}\log P_{\theta}(y|x_{w})+\textbf{odds}_{\theta}(y|x_{l})\cdot\nabla_{\theta}\log P_{\theta}(y|x_{l})\right)$ (20)
$=\nabla_{\theta}\log P_{\theta}(y|x_{w})(1+\textbf{odds}_{\theta}(y|x_{w}))-\nabla_{\theta}\log P_{\theta}(y|x_{l})(1+\textbf{odds}_{\theta}(y|x_{l}))$ (21)
The final equation is:
$\begin{split}\nabla_{\theta}\mathcal{L}_{I-OR}&=\left(1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}\right)^{-1}\cdot(\nabla_{\theta}\log P_{\theta}(y|x_{w})(1+\textbf{odds}_{\theta}(y|x_{w}))-\\ &\nabla_{\theta}\log P_{\theta}(y|x_{l})(1+\textbf{odds}_{\theta}(y|x_{l})))\end{split}$ (22)
$=\frac{1+\textbf{odds}_{\theta}(y|x_{w})}{1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}}\cdot\nabla_{\theta}\log P_{\theta}(y|x_{w})-\frac{1+\textbf{odds}_{\theta}(y|x_{l})}{1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}}\cdot\nabla_{\theta}\log P_{\theta}(y|x_{l})$ (23)
$=\left(1+\frac{\textbf{odds}_{\theta}(y|x_{w})}{\textbf{odds}_{\theta}(y|x_{l})}\right)^{-1}\cdot\left(\frac{\nabla_{\theta}\log P_{\theta}(y|x_{w})}{1-P_{\theta}(y|x_{w})}-\frac{\nabla_{\theta}\log P_{\theta}(y|x_{l})}{1-P_{\theta}(y|x_{l})}\right)$ (24)
Appendix E Preference Prompting
In this evaluation, we provide the model with the gold response $y$ and both instructions $x_{w}$ and $x_{l}$. We then prompt the model to choose the instruction most relevant to the gold text, following Bai et al. (2022) and Lee et al. (2023). The model should output ‘1’ if the first instruction generates the text and ‘2’ otherwise (Table 14). Next, we compare the log probabilities of the model outputting ‘1’ and ‘2’. If the log probability for ‘1’ is higher, we assume the model prefers whichever instruction came first in the prompt. The performance metric is determined by how often the model prefers the correct instruction, regardless of the order in which the correct instruction is presented. We experiment with Mistral-7b-Instruct-v0.2, Suri-I-ORPO, Suri-SFT, Mixtral-8x7b-Instruct-v0.1, Llama-3-7b-Instruct. All experiments use the Huggingface implementation with greedy decoding.
Prompt: p(preference|prompt) evaluation
You are an expert instruction rater. You will be given a text and two instructions, one of which is used to generate the text. Read through the text carefully, then determine which of the two instructions was used to generate the text. Answer only with "1" if the first instruction is correct, or "2" if the second instruction is correct. DO NOT give any reasoning.
### Text:
{text}
### First Instruction:
{ins1}
### Second Instruction:
{ins2}
Which instruction is correct? Answer only with "1" if the first instruction is correct, or "2" if the second instruction is correct. DO NOT give any reasoning.
Your response:
Table 14: Prompt used in the p(preference|prompt) evaluation. The {text} placeholder is replaced with gold responses, while the placeholders {ins1} and {ins2} are replaced with the correct and corrupted instructions, respectively. To mitigate any potential ordering bias, the order of the correct and corrupted instructions is shuffled. We will consider a response correct only if the model chooses the correct instruction, regardless of the ordering.
We observe that all models suffer from “first instruction bias", where the model always outputs the first instruction as the correct instruction, regardless of whether that instruction is actually $x_{w}$ or not.
Appendix F Human Evaluation
F.1 Recruitment
We recruit human annotators, all of whom are fluent in English, from Upwork (https://www.upwork.com) for our human evaluation. Each task is assigned to two annotators, except for Instruction Validation, which involves three annotators. Annotators are compensated at a rate of \$16 per hour and generally work an average of 12 hours per task. All annotators have signed consent forms, and our study has been approved by our institutional review boards (IRB).
F.2 Annotation
Figure 6 shows the LabelStudio interface for annotating instruction validity/constraint satisfaction. Figure 7 features the interface for comparing text generations based on how they satisfy a given constraint. Annotators note that the interfaces are user-friendly.
Figure 6: LabelStudio interface for annotating the validity of instructions. Annotators begin by carefully reading through the provided constraint and highlighting all the relevant text spans in the response supporting the constraint specified in the instruction. They then indicate whether the highlighted text satisfies the given constraint in the follow-up question.
Figure 7: LabelStudio interface for comparing generated text. Annotators begin by carefully reading through the provided constraint. They then highlight all the relevant text spans in the response that support the constraint specified in the instruction. After that, annotators answer questions on the informativeness, enjoyability, and coherence of the provided texts. We shuffle the generations in each task to prevent bias.
F.3 Annotator agreement in the instruction validity and constraint satisfaction evaluation
F.3.1 Power Analysis
We conduct a power analysis (Card et al., 2020) on our human evaluation data to estimate the number of text generations needed to achieve a power of 0.80 with a significance cutoff of 0.05 for our evaluation.
  1. 1.
    Evaluation 1 - Annotators rate text on constraint satisfaction (yes, no, partially): For this task, we estimated the effect size using Cohen’s w and employed a Chi-square goodness of fit test to determine the necessary number of text samples. We found that only 14.99 samples are needed to achieve a power of 0.80. In addition, with our current sample size of 30, we achieve a power of 0.9820, which exceeds the acceptable threshold of 0.80.
  2. 2.
    Evaluation 2 - Annotators indicate preference between text from model 1 and model 2: For this task, we estimated the effect size using Cohen’s h and used a z-test to calculate the required number of text samples. We determined that 23.17 samples are needed to reach a power of 0.80. With our sample size of 30, we achieve a power of 0.8902, which is above the acceptable threshold of 0.80.
While we evaluate 30 text samples for the first task, each text sample is evaluated with respect to approximately 10 constraints from the instructions, as we aim to account for constraints towards the end that are often missed by the models. Therefore, each annotator must evaluate a total of 321 (constraint, text) samples. This process is costly, with each annotator compensated \$200 and taking $\approx 15$ hours to complete the task.
F.3.2 Annotators’ Agreement
We note that Krippendorff’s Alpha remains low across evaluation tasks, suggesting little to no agreement among the annotators. We attribute this pattern to the fact that our generations are long ($\approx$4k words on average), making it hard for annotators to follow the narrative sometimes. Final statistics reported in the paper is averaged between the annotators.
Table 15 further shows disagreement types for the instruction validity and constraint satisfaction evaluation.
Types Krippendorff’s alpha Satisfied vs Partially Satisfied Partially Satisfied vs Not Satisfied Satisfied vs Not Satisfied
Instruction Validation (Section 2.3) 0.1 0.30 0.03 0.0
Constraint Satisfaction - Suri-SFT (Section 5) 0.0 0.52 0.24 0.24
Constraint Satisfaction - Suri-I-ORPO (Section 5) 0.2 0.60 0.34 0.06
I-ORPO vs SFT - Coherence (Section 5) 0.0 - - -
I-ORPO vs SFT - Informativeness (Section 5) 0.0 - - -
I-ORPO vs SFT - Enjoyability (Section 5) 0.1 - - -
Table 15: Types of disagreement among annotators in the instruction validation and constraint satisfaction tasks. Most disagreements arise over whether the text fully or partially satisfies the constraints.
Prompt: LLM evaluation
You will be given a text and its corresponding instruction, which contains the text’s main goal and a constraint. Determine whether the text satisfies the constraint (not the main goal). You should return your answer (Yes/No/Partially) along with your reasoning and a quote in the text that supports your reasoning (the quote should not contain any double quotation marks). Your answer should contain 3 fields: "answer", "reasoning", and "quote". DO NOT output anything else other than the response, which starts with "<<" and ending with ">>".
# Example 1: The text satisfies the constraint.
- Main goal: Write a first-person narrative describing a serene morning in a remote village.
- Constraint: You must not use the letter ’e’.
- Text: Dawn cracks with a yawn. On a hill, a hut sits, tranquil. Bright light climbs, casting gold on grass. In this calm morning, air is cool, birds sing softly. I stroll down paths, sipping hot cocoa, watching day start. Such is this dawn’s charm, lifting spirits, as world awakens.
- Your response: <<"answer": "Yes", "reasoning": "The text does not contain any ’e’, which satisfies the constraint.", "quote": "Dawn cracks with a yawn…">>
# Example 2: The text does not satisfy the constraint.
- Main goal: Compose a narrative that takes place entirely within the confines of a single, small room.
- Constraint: The story must not include any direct interaction or communication with other characters, whether through dialogue, notes, or any form of digital communication.
- Text: Sarah sat quietly in the corner of the small, dimly lit library room, surrounded by towering bookshelves filled with dusty volumes. Her focus was broken by a soft knock on the door. "Sarah, are you there?" her friend Emily’s voice called out gently from the other side. Sarah, startled yet relieved to hear a familiar voice, responded, "Yes, I’m here, Emily. Just give me a moment, I’ll open the door." They spent the next hour talking about the books Sarah had been reading and their plans for the weekend, making the small room feel a lot less lonely.
- Your response: <<"answer": "No", "reasoning": "The text includes a dialogue between Sarah and Emily, while the constraint specifies that the story must not include any direct interaction.", "quote": "’Sarah, are you there?’ her friend Emily’s voice called out gently from the other side. Sarah, startled yet relieved to hear a familiar voice, responded, ’Yes, I’m here, Emily. Just give me a moment, I’ll open the door.’">>
# Example 3: The text only satisfies part of the constraint.
- Main goal: Write a short story in which the protagonist meets an animal.
- Constraint: The walk should take place in a public space in a summer day.
- Text: As John strolled through the park one crisp autumn morning, he noticed the usual red and gold leaves blanketing the path. Today, however, a stray dog, thin and shivering, approached him. He hesitated, then offered his hand for the dog to sniff. It flinched at first, but soon warmed up to him. As they walked together, John wondered if he should take it home or find its owner.
- Your response: <<"answer": "Partially", "reasoning": "The text mentions that the character walks in a park, which satisifies the constraint that the setting is a public place. However, the walk takes place in an autumn morning, which violates the constraint that the walk takes place in a summer day", "quote": "As John strolled through the park one crisp autumn morning, he noticed the usual red and gold leaves blanketing the path…">>
# Instruction
## Main Goal
{goal}
## Constraint
{constraint}
# Text
{text}
DO NOT output anything else other than the response, which starts with "<<" and ending with ">>".
# Your response
Table 16: Prompt to evaluate whether a text follows a constraint or not. {goal}, {constraint}, and {text} are placeholders that will be replaced with actual content.
Appendix G Generations from Mistral-Instruct, Suri-SFT, and Suri-ORPO
We show generations from Mistral-Instruct, Suri-SFT, and Suri-ORPO in Table 17.
Constraint Suri-I-ORPO Suri-SFT Mistral-Instruct
Detail the onset of a serious, life-altering illness, including the initial diagnosis and the impact it had on the quality of life and daily activities of the person. Being diagnosed with MSA often means a rapid decline, and it didn’t take long for it to rob Elizabeth of her steady gait and coordination. There were numerous falls, steps hit, food dropped. There was depression, panic, and the wrenching witness to a vibrant, funny, independent woman being stripped of her abilities. Slowly, as she grew more debilitated, we made adjustments. She stopped driving, which was an excruciating blow to her. She then lost her ability to use her laptop. She struggled to get into the shower, and then used a shower chair. She used a wheelchair, and then an electric scooter. She became a master of calling an ambulance before her body decided to keep her upright. She was diagnosed with advanced pulmonary hypertension in July of 2013, immediately after heart failure, which was the result of an array of intestinal diseases and surgeries that dated back to 2001…For several years, she was able to continue attending to her business and loved entertaining friends and partying at home. However, as her illness advanced, it became impossible for her to live the life that she loved. She was no longer able to play her beloved music, do needle point, bake, cook or garden. She struggled to even go for walks, even though she had many appointments at local hospitals and clinics. She became primarily housebound, with only short trips to the grocery store or to see a pet warden. She was reduced to having nearly 24hrs a day of assistance, administration of a nebulizer, and hospice nurses intervening if her symptoms increased. But the seasons turned, and so did Amelia’s fate. One fateful autumn day, she received the news that would forever change her life: a diagnosis of a degenerative neurological condition. The prognosis was grim, a cruel reminder that even the strongest among us are mortal. Her once nimble hands could no longer hold a pen to write her poetry, her once steady legs could no longer carry her to tend to her garden. Her light began to dim, and the darkness of her illness cast a long, tragic shadow over her life.
Begin the narrative with a detailed description of the setting, emphasizing the time of day, the condition of the sea, and the immediate surroundings to establish a strong sense of place. The ship cut through the waves as evening fell. The sky was painted the colors of a summer sunset, deepening to pink, orange, purple, and green in the west as the sun sank toward the horizon. The air was warm, redolent with salt, and the ship rocked with the weight of the cargos that it carried in its hold. Ahead, the Pale Horse sped through the waves, its white sails snapping in the wind that billowed from the west. A flock of seabirds wheeling above the ship took no notice of its passage. Nearby, two smaller vessels followed, one with a low, curving bow that gave it an almost feminine grace in the breeze, the other a large merchant ship, weighed down with grain and other trade items. To the south east, the Pale Horse’s sister ship, the Red Lion, was already out of sight. The sun was going down, a pink glow still covering the horizon, but its fading light made the water look almost pitch black. A light surf broke against the stones of the harbor’s seawall. High on the shore, where the promontory formed a horseshoe shape, there was a small port, and beyond that the main wall of the city, the tower of the Castello Giustinian, and the cathedral could be seen against the blackening sky. Here, off-shore, there were only the dark blue waters of the Adriatic. There was no color, no life, and the sea stretched away for miles. The evening sun began its descent towards the western horizon, casting long shadows on the undulating waters of the River Tiber. The air, thick with the fragrant bouquet of jasmine and rose from the gardens of Palazzo Borgia, was tinged with a hint of chill as autumn’s tender tendrils crept closer to the heart of the Italian Peninsula.
Table 17: Example generations from Suri-I-ORPO, Suri-SFT, Mistral-Instruct. All three generations make a decent attempt at following the given constraint.
Appendix H Comparable training setup for Suri-SFT and Suri-I-ORPO
Here, we report the results for Suri-SFT-single, where we fine-tune Mistral-Instruct-7B-v0.2 using the same instruction setup as I-ORPO (including only one constraint in the instruction). Table 18 demonstrates that Suri-SFT-single underperforms Suri-I-ORPO in all aspects. While Suri-SFT-single does generate longer text compared to the baseline models, it exhibits more repetitions and lower ranking accuracy than both Suri-I-ORPO and Suri-SFT. Our qualitative analysis of 30 text samples generated by Suri-SFT-single shows that the generations contain more gibberish, satisfy fewer constraints, and are generally harder to read. These findings reinforce our original claim that I-ORPO outperforms SFT.
Metric Suri-SFT-Single Suri-I-ORPO
Average number of tokens in test set generation 4007 5102
Number of test set generations with 5-gram repetitions 32% 24%
Number of test set generations with 10-gram repetitions 4% 3%
Ranking accuracy [$(M,M)$, $(M,M/2)$, $(M,1)$, $(M/2,M/2)$, $(1,1)$] [99, 94, 65, 96, 76] [100, 100, 98, 100, 98]
Table 18: Comparison of Suri-SFT-Single and Suri-I-ORPO performance metrics
Appendix I Fine-tuning on Suri does not significantly degrade performance in short-form instruction following tasks
We measure the performance of Suri-I-ORPO and Suri-SFT on popular benchmarks using lm-evaluation-harness (Gao et al., 2024). Our findings indicate that our fine-tuned models do not significantly degrade the performance of the baseline instruct model (Table 19). In fact, they slightly improve performance on most benchmarks, with the exceptions of HellaSwag, WinoGrande, and Arc-challenge.
Model MMLU HellaSwag WinoGrande PIQA Arc-e Arc-c GSM8K GPQA
I-ORPO 62.81% 80.31% 72.69% 80.63% 81.99% 53.67% 42.23% 31.25%
SFT 57.79% 80.21% 71.82% 81.18% 82.03% 53.07% 41.09% 30.25%
Mistral-7b-instruct-v2 (baseline) 58.73% 83.67% 73.48% 80.25% 81.36% 54.44% 42.08% 26.34%
Table 19: Comparison of model performance across various tasks
Appendix J GPT-4o-mini’s performance on Suri
Model GPT-4o-mini Suri-I-ORPO
Average number of tokens in test set generation 1134 5102
Number of test set generations with 5-gram repetitions 6% 24%
Number of test set generations with 10-gram repetitions 1% 3%
Table 20: Comparison between GPT-4o-mini and Suri-I-ORPO on test set generations.
We note that although GPT-4o-mini produces less repetitive text, it can generate only an average of 1,134 tokens, which is still lower than Mixtral-8x7B-Instruct (Table 20). The higher repetition rate in the fine-tuned models may simply be due to these models generating longer text. Upon analyzing 30 generation samples from GPT-4o-mini, we observe that while the model can satisfy the constraints, it still suffers from formulaic generation and unnatural incorporation of those constraints.